OpenStudy (anonymous):

how to solve 2cos^2 x- 3cosx +1=0 on the interval {0,2pi}

OpenStudy (solomonzelman):

let `cos(x) = a` then solve the quadratic....

OpenStudy (anonymous):

I did it and,i got cos x=1 and cos x = 1/2 but I need to know on the interval {0,2pi} what is the solution

OpenStudy (anonymous):

You are correct so far.

OpenStudy (anonymous):

Now you have \(\cos x = 1\) or \(\cos x = \dfrac{1}{2} \)

OpenStudy (anonymous):

Let's deal with \(\cos x = 1\) first. Here is a unit circle. Which angle measure from 0 to \(2\pi\) has a cosine of 1?